# Simple and efficient methods for local structural analysis in polydisperse hard disk systems

### v1

@@ II.2.4 Detecting confirmed disks for constructing free volume
    Employ extended 2D-SANN to detect neighboring disks within the $n$th shells (e.g., $n=2$) for tagged disk $i$. Include a sufficient number of candidate disks $\{j\}$ to consider the free volume of $i$.
@@
    Follow the same procedures as in steps 3–6 of NELF-A by 2D-SANNex.
@@ III.1 Numerical settings
We demonstrate the differences between the methods with a typical equilibrated configuration in systems of both mono- and bidisperse elastic hard disks in two dimensions. In the investigation of hard sphere/disk models, the development of computer simulation methods, namely, Monte Carlo (MC)[36] and molecular dynamics (MD)[37], represented a milestone, enabling researchers to unveil the mysteries of the phase transitions deriving from the entropic force[38, 39](the so-called Alder transition). In hard disk systems[40], the controversy regarding the type of phase transitions has now been settled by massive calculations using three modern methods[41, 42]. Hard disk systems are subject to two phase transitions, between fluid, hexatic, and solid phases, at around $\nu\sim 0.700$ and $0.716$, respectively. The free volume of a hard sphere is an essential quantity for understanding phenomena such as phase transitions and the coarsening process of phase separation (entropic force) in binary mixtures approaching the glass transition[2, 8].
@@
Table 1: Comparison of NN calculations between the fixed-distance cutoff, 2D-SANN, and 2D-SANNex algorithms and of free volume calculations between (a) NELF-A by 2D-SANNex, (b) NELF-A by cutoff, and (c) NELF-A by first finding all intersections for mono- and bidisperse systems.
@@ III.2 Efficiency of NN and free volume estimators
In addition to NN calculations, we also examined the efficiencies of the various NELF-A methods described in Sec. II.2.4, as presented in Table 1. In NELF-A by 2D-SANNex was used, and NNs were calculated up to the second shells. For monodisperse systems, NELF-A by first finding all intersections was the fastest, completing the calculation in about half the time taken by the other two methods. Conversely, for bidisperse systems, NELF-A by cutoff demonstrated the best efficiency. NELF-A by first finding all intersections was approximately two orders of magnitude slower than the other two methods for bidisperse systems, and it was slower than for monodisperse systems. We attribute this to the strong influence of particle dispersions on NELF-A and the consequent need to calculate and memorize the intersections of all particles in the system for every dispersion. As discussed in Sec. II.2.4, considering factors such as particle dispersions and the computer’s memory capacity, different algorithms are necessary for efficient computing, depending on the situation. In Sec. III.4, NELF-A by first finding all intersections is used for monodisperse systems, while NELF-A by cutoff is used for bidisperse systems.
@@ III.3.1 Local distributions of the number of NNs
We then focused on the differences in the more localized spatial distributions of NNs between methods. Focusing on the large disk located at the center of each enlarged view in Fig. 12, $7$NNs are identified by the Voronoi construction, while the other two methods identify $5$NNs. A small part of the edge of a polygon, between the large and small disks (located on the upper right of the large disk), is shared and recognized as an NN, i.e., as topologically connected. However, these disks are located relatively far from the other NNs; i.e., the metric distance is large. The instability in the number of neighbors is expected to be caused by thermal fluctuations, which results in the detection of a relatively large number of NNs. This is identified as one of the disadvantages of the Voronoi construction.
@@ III.3.2 Probability distribution of the number of NNs
The probability distribution functions (PDFs) of the number of NNs in bidisperse hard disk systems at $\nu=0.720$ were calculated using $300$ frame-equilibrated configurations at $(N,\nu)=(64^{2},0.720)$, as shown in Fig. 13. Here, the number of NNs is estimated by the fixed-distance cutoff (C), the modified cutoff (mC), Voronoi construction (V), and 2D-SANN (S), respectively. A notable difference with 2D-SANN compared with the other methods is that $5$NNs has the highest probability, whereas $6$NNs is most probable according to the other methods. This trend remains consistent at $\nu=0.760$ and $0.780$. Additionally, with the modified cutoff method, where multiple cutoff distances corresponding to each particle pair are introduced, it is observed that the numbers of $5$NNs and $7$NNs are higher compared with the fixed-distance cutoff. Furthermore, the Voronoi construction identifies more $7$NNs and fewer $5$NNs than the fixed-distance cutoff. This observation that a relatively large number of NNs are detected is expected, since disks located at large metric distances are topologically shared with the short edges.
@@
Figure 14 shows a comparison of the PDFs of the number of second neighbors according to the extended cutoff method[23] (eC) and extended 2D-SANN (eS) (see Sec. II.1.4 and Figs. 7 and 6). With the extended cutoff, from Fig. 2, the cutoff radii for the first and second shells are set as $r_{\mathrm{c}}^{\ast}=r_{\mathrm{c}}/(2\sigma_{\mathrm{e}})=1.49$ and $2.52$, respectively. Extended 2D-SANN is found to give a smaller value for the most frequent number of second neighbors compared with the extended cutoff, and this trend is similar to the results for the (first) NNs. Additionally, as the particle packing fraction increases to $\nu=0.780$, in the extended cutoff method, $12$NNs becomes more dominant, whereas, with extended 2D-SANN, the probabilities of $11$NNs and $12$NNs become almost the same, indicating that both $11$NNs and $12$NNs become dominant as the packing fraction increases.
@@ IV Concluding Remarks
In this paper, we have discussed different the methods for nearest neighbor and free volume estimation, including the higher-order treatments of polydisperse hard disk systems. We have clarified the difficulties that arise with conventional methods using cutoffs and Voronoi constructions in local microscopic environment analyses of molecular configurations. We have confirmed that the solid-angle-based nearest-neighbor (SANN) algorithm is able to overcome the difficulties encountered by conventional methods, even for polydisperse hard disk systems. We have also presented an explicit definition for evaluating higher-order neighbors. The approaches proposed here for the determination of nearest neighbors and the free volume provide crucial information to elucidate the dynamics of slow relaxation and hopping motion, especially in densely packed systems. We have proposed a simple, efficient, and precise method for categorizing neighbors for enclosing the local free area (NELF-A) in the case of dense polydisperse hard disk systems.
In future extensions of this work, we aim to investigate its applicability to (i) finding inherent structure in dense polydisperse hard disk systems, (ii) systems with various boundary conditions and (iii) systems in three and more dimensions, which remain an open problem.

### v2

@@ II.2.4 Detecting confirmed disks for constructing free volume
    Employ extended 2D-SANNex to detect neighboring disks within the $n$th shells (e.g., $n=2$) for tagged disk $i$. Include a sufficient number of candidate disks $\{j\}$ to consider the free volume of $i$.
@@
    Follow the same procedures as in steps 3–6 of NELF-A by extended 2D-SANNex.
@@ III.1 Numerical settings
We demonstrate the differences between the methods with a typical equilibrated configuration in systems of both mono- and bidisperse elastic hard disks in two dimensions. In the investigation of hard sphere/disk models, the development of computer simulation methods, namely, Monte Carlo (MC)[36] and molecular dynamics (MD)[37], represented a milestone, enabling researchers to unveil the mysteries of the phase transitions deriving from the entropic force[38, 39](the so-called Alder transition). In hard disk systems[40], the controversy regarding the type of phase transitions has now been settled by massive calculations using three modern methods[41, 42]. Hard disk systems are subject to two phase transitions, between fluid, coexisting, hexatic, and solid phases, at around $\nu\sim 0.700$, $0.716$, and $0.720$, respectively. The free volume of a hard sphere is an essential quantity for understanding phenomena such as phase transitions and the coarsening process of phase separation (entropic force) in binary mixtures approaching the glass transition[2, 8].
@@
Table 1: Comparison of NN calculations between the fixed-distance cutoff, 2D-SANN, and 2D-SANNex algorithms and of free volume calculations between (a) NELF-A by extended 2D-SANNex, (b) NELF-A by cutoff, and (c) NELF-A by first finding all intersections for mono- and bidisperse systems.
@@ III.2 Efficiency of NN and free volume estimators
In addition to NN calculations, we also examined the efficiencies of the various NELF-A methods described in Sec. II.2.4, as presented in Table 1. In the case of NELF-A by extended 2D SANNex, neighboring disks were calculated up to the second shells. For monodisperse systems, NELF-A by first finding all intersections was the fastest, completing the calculation in about half the time taken by the other two methods. Conversely, for bidisperse systems, NELF-A by cutoff demonstrated the best efficiency. NELF-A by first finding all intersections was approximately two orders of magnitude slower than the other two methods for bidisperse systems, and it was slower than for monodisperse systems. We attribute this to the strong influence of particle dispersions on NELF-A and the consequent need to calculate and memorize the intersections of all particles in the system for every dispersion. As discussed in Sec. II.2.4, considering factors such as particle dispersions and the computer’s memory capacity, different algorithms are necessary for efficient computing, depending on the situation. In Sec. III.4, NELF-A by first finding all intersections is used for monodisperse systems, while NELF-A by cutoff is used for bidisperse systems.
@@ III.3.1 Local distributions of the number of NNs
We then focused on the differences in the more localized spatial distributions of NNs between methods. Focusing on the large disk located at the center of each enlarged view in Fig. 12, $7$NNs are identified by the Voronoi construction, while the other two methods are identified by $5$NNs. A small part of the edge of a polygon, between the large and small disks (located on the upper right of the large disk), is shared and recognized as an NN, i.e., as topologically connected. However, these disks are located relatively far from the other NNs; i.e., the metric distance is large. The instability in the number of neighbors is expected to be caused by thermal fluctuations, which results in the detection of a relatively large number of NNs. This is identified as one of the disadvantages of the Voronoi construction.
@@ III.3.2 Probability distribution of the number of NNs
The probability distribution functions (PDFs) of the number of NNs in bidisperse hard disk systems at $\nu=0.720$ were calculated using $300$ frame-equilibrated configurations at $(N,\nu)=(64^{2},0.720)$, as shown in Fig. 13. Here, the number of NNs is estimated by the fixed-distance cutoff (C), the modified cutoff (mC), Voronoi construction (V), and 2D-SANN (S), respectively. A notable difference with 2D-SANN compared with the other methods is the case of $5$NNs has the highest probability, whereas the case of $6$NNs is most probable according to the other methods. This trend remains consistent at $\nu=0.760$ and $0.780$. Additionally, with the modified cutoff method, where multiple cutoff distances corresponding to each particle pair are introduced, it is observed that the numbers of $5$NNs and $7$NNs are higher compared with the fixed-distance cutoff. Furthermore, the Voronoi construction identifies more $7$NNs and fewer $5$NNs than the fixed-distance cutoff. This observation that a relatively large number of NNs are detected is expected, since disks located at large metric distances are topologically shared with the short edges.
@@
Figure 14 shows a comparison of the PDFs of the number of second neighbors according to the extended cutoff method[23] (eC) and extended 2D-SANN (eS) (see Sec. II.1.4 and Figs. 6 and 7). With the extended cutoff, from Fig. 2, the cutoff radii for the first and second shells are set as $r_{\mathrm{c}}^{\ast}=r_{\mathrm{c}}/(2\sigma_{\mathrm{e}})=1.49$ and $2.52$, respectively. Extended 2D-SANN is found to give a smaller value for the most frequent number of second neighbors compared with the extended cutoff, and this trend is similar to the results for the (first) NNs. Additionally, as the particle packing fraction increases to $\nu=0.780$, in the extended cutoff method, $12$NNs becomes more dominant, whereas, with extended 2D-SANN, the probabilities of $11$NNs and $12$NNs become almost the same, indicating that both $11$NNs and $12$NNs become dominant as the packing fraction increases.
@@ IV Concluding Remarks
In this paper, we have discussed different the methods for nearest neighbor and free volume estimation, including the higher-order treatments of polydisperse hard disk systems. We have clarified the difficulties that arise with conventional methods using cutoffs and Voronoi constructions in local microscopic environment analyses of molecular configurations. We have confirmed that the solid-angle-based nearest-neighbor (SANN) algorithm is able to overcome the difficulties encountered by conventional methods, even for polydisperse hard disk systems. We have also presented an explicit definition for evaluating higher-order neighbors, and have proposed a simple, efficient, and precise method for categorizing neighbors for enclosing the local free area (NELF-A) in the case of dense polydisperse hard disk systems. The approaches proposed here for the determination of nearest neighbors and the free volume provide crucial information to elucidate the dynamics of slow relaxation and hopping motion, especially in densely packed systems.
In the future, we aim to further explore the applications of NELF-A by focusing on two main objectives: (i) developing a methodology for deriving inherent structures in hard disk systems using NELF-A, and (ii) extending the applicability of NELF-A to systems with various boundary conditions and three or more dimensions.